Grade 11Trigonometry

what is trigonometric expansions?

nithin kumar ashtekar
16 Years agoGrade 11

1 Answer

Badiuddin askIITians.ismu Expert
ApprovedApproved Tutor Answer16 Years ago

Dear nithin

Representation of power of trignometric ratio in terms trignometric ratio of multiple angel is known as trignometric expansions

 

for example.

cos6q = 1/32{cos 6q+6cos 4q+15cos 2q+10 }

or

cos7q = 1/64{cos 7q+7cos 5q+21cos 3q+35cos q  }

 

these expantion can bederived by puting cosx =(eix+e-ix)/2

  and expand it with the hepl of binomial .

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin